# Heisenberg homology of ribbon graphs

### v2

@@ 1 Heisenberg homology of surface configurations
The surface braid group is then defined as $\mathbb{B}_{n}(\Sigma_{g,m})=\pi_{1}(\mathcal{C}_{n}(\Sigma_{g,m}),*)$. A presentation for this group was first obtained by G. P. Scott [4] and revisited by Gonzáles-Meneses [5], Bellingeri [6]. We fix based loops, $\alpha_{1},\dots,\alpha_{g},\beta_{1},\dots,\beta_{g},\gamma_{1},\dots,\gamma_{m-1}$, as depicted in Figure 1. The base point *$*_{1}$ belongs to the base configuration ***. We will use the same notation $\alpha_{r}$, $\beta_{s}$, $\gamma_{t}$, for the corresponding braids where only the first point is moving.
@@
$\begin{cases}\,\text{({BR1}) }\,[\sigma_{i},\sigma_{j}]=1&\text{for }\lvert i-j\rvert\geqslant 2,\\ \,\text{({BR2}) }\,\sigma_{i}\sigma_{j}\sigma_{i}=\sigma_{j}\sigma_{i}\sigma_{j}&\text{for }\lvert i-j\rvert=1,\\ \,\text{({CR1}) }\,[\zeta,\sigma_{i}]=1&\text{for }i>1\text{ and all }\zeta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\\ \,\text{({CR2}) }\,[\zeta,\sigma_{1}\zeta\sigma_{1}]=1&\text{for all }\zeta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\\ \,\text{({CR3}) }\,[\zeta,\sigma^{-1}_{1}\eta\sigma_{1}]=1&\text{for all }\zeta\neq\eta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\text{ with}\\ &\{\zeta,\eta\}\neq\{\alpha_{r},\beta_{r}\},\\ \,\text{({SCR}) }\,\sigma_{1}\beta_{r}\sigma_{1}\alpha_{r}\sigma_{1}=\alpha_{r}\sigma_{1}\beta_{r}&\text{for all }r.\end{cases}$ (1)
@@ 3 A complex for Borel-Moore homology of graph configurations
The Borel-Moore homology module $H_{n}^{BM}(\mathcal{C}_{n}(\Sigma_{g,m})),\mathcal{C}_{n}(\Sigma_{g,m},A);W)$ is isomorphic to the direct sum of $\left(\begin{array}[]{c}2g+m+n-2\\ n\\ \end{array}\right)$ copies of $W$. Furthermore, $H_{*}^{BM}(\mathcal{C}_{n}(\Sigma_{g,m}),\mathcal{C}_{n}(\Sigma_{g,m},A);W)$ vanishes for *$*\neq n$.
@@ 4.1 Action on the Heisenberg group
[…]

### v3

@@ 1 Heisenberg homology of surface configurations
The surface braid group is then defined as $\mathbb{B}_{n}(\Sigma_{g,m})=\pi_{1}(\mathcal{C}_{n}(\Sigma_{g,m}),*)$. A presentation for this group was first obtained by G. P. Scott [4] and revisited by Gonzáles-Meneses [5], Bellingeri [6]. We fix based loops, $\alpha_{1},\dots,\alpha_{g},\beta_{1},\dots,\beta_{g},\gamma_{1},\dots,\gamma_{m-1}$, as depicted in Figure 1. The base point $*_{1}$ belongs to the base configuration *. We will use the same notation $\alpha_{r}$, $\beta_{s}$, $\gamma_{t}$, for the corresponding braids where only the first point is moving.
@@
$\begin{cases}\,\text{({BR1}) }\,[\sigma_{i},\sigma_{j}]=1&\text{for }\lvert i-j\rvert\geqslant 2,\\ \,\text{({BR2}) }\,\sigma_{i}\sigma_{j}\sigma_{i}=\sigma_{j}\sigma_{i}\sigma_{j}&\text{for }\lvert i-j\rvert=1,\\ \,\text{({CR1}) }\,[\zeta,\sigma_{i}]=1&\text{for }i>1\text{ and all }\zeta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\\ \,\text{({CR2}) }\,[\zeta,\sigma_{1}\zeta\sigma_{1}]=1&\text{for all }\zeta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\\ \,\text{({CR3}) }\,[\zeta,\sigma^{-1}_{1}\eta\sigma_{1}]=1&\text{for all }\zeta\neq\eta\text{ among the }\alpha_{r},\beta_{s},\gamma_{t},\text{ with}\\ &\{\zeta,\eta\}\neq\{\alpha_{r},\beta_{r}\},\text{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} $\zeta$ on the left of $\eta$},\\ \,\text{({SCR}) }\,\sigma_{1}\beta_{r}\sigma_{1}\alpha_{r}\sigma_{1}=\alpha_{r}\sigma_{1}\beta_{r}&\text{for all }r.\end{cases}$ (1)
@@ 3 A complex for Borel-Moore homology of graph configurations
The Borel-Moore homology module $H_{n}^{BM}(\mathcal{C}_{n}(\Sigma_{g,m})),\mathcal{C}_{n}(\Sigma_{g,m},A);W)$ is isomorphic to the direct sum of $\left(\begin{array}[]{c}2g+m+n-2\\ n\\ \end{array}\right)$ copies of $W$. Furthermore, $H_{*}^{BM}(\mathcal{C}_{n}(\Sigma_{g,m}),\mathcal{C}_{n}(\Sigma_{g,m},A);W)$ vanishes for $*\neq n$.
@@ 4.1 Action on the Heisenberg group
${\mathbb{B}_{n}(\Sigma)}$${\mathbb{B}_{n}(\Sigma)}$${\mathcal{H}(\Sigma)}$${\mathcal{H}(\Sigma)}$$\scriptstyle{\phi}$$\scriptstyle{f_{\mathbb{B}_{n}(\Sigma)}}$$\scriptstyle{\phi}$$\scriptstyle{f_{\mathcal{H}}}$ (15)